# Pluto and Charon: A Time and Memory Efficient Collaborative Edge AI Framework for Personal LLMs Fine-Tuning

### v1

@@ 6.2. End-to-end Performance
(b) Peak memory consumption of LLM weights per device across the edge cluster.
Figure 9. Comparison of different fine-tuning techniques.
@@
Figure 10. Device grouping results of PAC’s hybrid parallelism for experiments in Figure 9. "N" indicates Jetson Nano.
6.3. Significance of Parallel Adapters at the Edge
We conducted experiments to assess the time and memory efficiency of Parallel Adapters at the edge. In this section, we perform data parallelism for Parallel Adapters with activation cache across 8 devices and hybrid parallelism for other fine-tuning techniques without 1F1B micro-batch scheduling. "Activations" contain the intermediate results and optimizer states. Figure 8 illustrates that Parallel Adapters outperform other fine-tuning techniques regarding both time and memory efficiency.
Parallel Adapters markedly reduce per-sample training time. Figure 8(a) presents the average sample training time across different fine-tuning techniques. Without activation cache, Parallel Adapters can reduce the average sample training time by $31.94\%$ to $56.24\%$ compared to baseline methods, primarily owing to a substantial decrease in backward propagation overhead. Both Adapters and LoRA incorporate trainable structures into the backbone model, thus necessitating backpropagation across the entire backbone model for gradient computation of these parameters. Consequently, regarding backward time, Adapters and LoRA can only achieve approximately a $49\%$ reduction compared to full fine-tuning. In contrast, backpropagation through the backbone model is unnecessary with Parallel Adapters, leading to a more substantial reduction in backward time, nearly $92\%$ compared to full fine-tuning. Moreover, Parallel Adapters With activation cache mechanism can further decrease the average sample training time up to $96.39\%$. These results demonstrate the substantial reduction in training time achieved by Parallel Adapters.
Parallel Adapters yield a substantial reduction in memory usage. Figure 8(b) depicts the breakdown of the memory footprint for different fine-tuning techniques. We report the peak memory consumption per device across edge clusters. Without activation cache, Parallel Adapters can reduce memory usage by $25.27\%$ to $60.49\%$. Adapters and LoRA demonstrate parameter efficiency but do not exhibit significant memory efficiency. While these techniques notably decrease the memory footprint of gradients by reducing the number of trainable parameters, the memory usage associated with activations remains considerable. However, intermediate activations often become the primary memory bottleneck during training, especially with larger batch sizes. For PEFT techniques such as Adapters and LoRA, activation memory can be reduced by up to $28.15\%$ compared to full fine-tuning across the three models. In contrast, Parallel Adapters achieve a more significant reduction, reaching up to $58.87\%$. With activation cache, Parallel Adapters can decrease the peak memory footprint from $74.57\%$ to $88.16\%$ compared to baselines. This is because it’s sufficient to store only the lightweight Parallel Adapters, eliminating the need to host the entire LLM backbone in memory.
6.4. Analysis of Collaborative Edge Fine-Tuning
We perform an ablation study to understand the contribution of hybrid parallelism and activation cache in our system design.
@@ 6.4. Analysis of Collaborative Edge Fine-Tuning
Comparasion PAC with EDDL and Eco-FL. To explore the scalability advantages of PAC’s hybrid parallelism over Eco-FL’s pipeline parallelism and EDDL’s data parallelism, we compared the throughput of these methods when training collaboratively across 2 to 8 edge devices. The batch size was consistent with the number of devices, and the sequence length of each sample was fixed at 128. We implement Eco-FL and EEDL using the Parallel Adapters technique to ensure a fair comparison. Note that none of the three methods utilizes activation cache.
Figure 9(b) illustrates the maximum per device memory footprint of model weights across edge cluster. For EDDL, each device must host a complete LLM, preventing the reduction of the parameters’ memory footprint through scaling up the number of devices. Therefore, as shown in Figure 9(a), the EDDL method exhibits OOM errors with both the BART-Large and T5-Large models. Conversely, PAC and Eco-FL utilize pipeline parallelism, partitioning the model into multiple stages with each handled by different devices. This approach allows for scaling the number of devices to reduce the peak memory footprint. PAC’s hybrid parallelism offers a broader search space for parallel strategies compared to Eco-FL’s pipeline parallelism. Our planning algorithm for PAC is capable of identifying more efficient hybrid parallel configurations within memory constraints, enhancing resource utilization. Although PAC may incur higher memory overhead in some instances, it achieves greater system throughput. Specifically, when compared to Eco-FL, PAC exhibits an increase in throughput from $39.50\%$ to $84.79\%$.
@@
Figure 11. Fine-tuning time with PAC. Time without activation cache is represented by bars. The corresponding reduction in time achieved utilizing activation cache is represented by shaded areas. Dataset: MRPC.
@@
To more clearly illustrate the parallel strategies adopted by PAC, we present the device grouping configurations for PAC across various LLMs and numbers of devices in Figure 10. On the left, a table displays all the grouping results across three models. On the right, an instance is shown where a model is divided into three stages, with two devices handling the first stage to perform data parallelism. Specifically, when fine-tuning BART-Large with eight devices, EDDL encounters OOM issues because a single Jetson Nano cannot accommodate a complete BART-Large. Eco-FL addresses this problem by dividing the model into eight stages and employing straight pipeline parallelism for training. On the contrary, our PAC approach divides BART-Large into two stage models, with each stage replicated across four devices. This configuration significantly reduces the number of stages in the pipeline, thereby minimizing inter-stage data dependencies and communication latency, which in turn enhances the pipeline’s concurrent efficiency. These results demonstrate that our hybrid parallel approach offers a larger search space for parallel configurations, providing enhanced scalability and robustness across varying numbers of devices and workloads.
@@
Comparison of PAC with and without activation cache. We further investigated how our activation cache mechanism benefits the required fine-tuning latency. By leveraging activation cache, the fine-tuning latency per epoch can decrease up to $79.51\%$. Figure 11 shows the fine-tuning latency reduction as the number of epochs increases with the use of the activation cache mechanism. We can observe that by leveraging the activation cache, fine-tuning latency is reduced by $26\%$ to $71\%$. Moreover, this reduction in latency increases with the number of epochs. For example, with the T5-Large model, training for two epochs reduces latency by $39\%$, whereas training for ten epochs increases the reduction to $71\%$. This reduction can be attributed to the fact that the Parallel Adapters constitute a lightweight, independent network, resulting in a significant decrease in training cost compared to the LLM backbone. We can bypass both the forward and backward passes through the LLM backbone since the necessary activations are already cached.
7. Related Work
@@
Parameter-Efficient Fine-Tuning for LLM. Prompt tuning (Lester et al., 2021) proposes to prepend the model input embeddings with a trainable tensor. Adapters tuning (Houlsby et al., 2019) adds domain-specific layers after attention and FFN layers in transformer. LoRA (Hu et al., 2021) decomposes the parameter update for a weight matrix into two trainable low-rank matrices. To further reduce the memory overhead, pioneering studies explore fine-tuning techniques that obviate the need for backpropagation through the backbone model. Y-tuning (Liu et al., 2024) learns additional task-specific label representations, which are integrated with the output of the backbone model to circumvent backpropagation. LST (Sung et al., 2022) involves the use of pruned lightweight transformer structures from the backbone as a side network. E3VA (Yin et al., 2023) extends the concept of the side network into the realm of computer vision.
@@
On-device DNN Fine-Tuning. POET (Patil et al., 2022) achieves the finetuning of a BERT model on embedded devices, optimizing for both training speed and energy consumption. Lin et al. (Lin et al., 2022) enable training directly on devices with a minimal memory requirement of only 256KB. Sage and Melon (Gim and Ko, 2022; Wang et al., 2022) implement hybrid memory management and conservation strategies, including operator fusion and the use of a dedicated memory pool, to mitigate memory limitations. Additionally, Mandheling (Xu et al., 2022) incorporates mixed-precision training along with DSP offloading to enhance the speed of learning.
@@
Collaborative Edge Computing for DNN Fine-Tuning. Federated Learning (FL) has been a promising paradigm in distributed machine learning that enables in-situ model fine-tuning. FwdLLM (Xu et al., 2024) designs a backpropagation-free fine-tuning FL protocol to enhance efficiency. AdaFL (Cai et al., 2023) proposes an FL framework for fine-tuning LLMs that features adaptable depth and width in its adapters modules. Breaking through the conventional paradigm of FL, Ye et al. (Ye et al., 2022; Zeng et al., 2024) devise a pipeline parallel architecture that facilitates the collaborative fine-tuning of DNNs across multiple edge devices. EDDL (Hao and Zhang, 2021) adopts data parallelism training across embedded devices in a local area network. Asteroid (Ye et al., 2024b) also employs HPP across multiple edge devices for DNN training, but it does not specifically address the parameter-efficient fine-tuning of LLMs.
8. Conclusion
@@
This paper proposes PAC, a time and memory efficient collaborative edge AI framework for personal LLMs fine-tuning. PAC breaks the resource wall of personal LLMs fine-tuning with a sophisticated algorithm-system co-design, achieving a acceleration of 8.64×8.64$\times$8.64 × and $88.16\%$ memory reduction compared to state-of-the-art methods.

### v2

@@ VI-E Analyzing the Impact of Weight Initialization on Parallel Adapters
(b) Peak memory consumption of LLM weights per device.
Figure 16: Comparison of different collaborative edge computing architecture.
@@
Figure 17: Device grouping results of PAC+’s hybrid parallelism for experiments in Figure 16. "N" indicates Jetson Nano.
VI-F Analysis of Fine-Tuning with Quantized LLMs
(1) Memory Footprint Analysis. We compare the memory footprint during the fine-tuning process of LLMs with different number of parameters, using full-parameter fine-tuning, LoRA, Adapter, and Parallel Adapters algorithms with various quantization precisions. We obtained a series of T5 models with varying parameter sizes by adjusting parameters such as hidden size, number of layers, and number of attention heads. Experimental results show that Parallel Adapters with lower quantization precisions, such as INT4 and INT8, can significantly reduce the memory footprint during fine-tuning. Specifically, compared to full-parameter fine-tuning, the Parallel Adapter algorithm with INT4 precision reduces memory footprint by up to 88%, and compared to existing PEFT algorithms, it can reduce memory footprint by up to 73%.
(2) Analysis of Fine-Tuning Performance with Quantized LLMs. Building on Table VI, we further evaluate and report the final fine-tuned performance of Parallel Adapters with different quantization precisions across four datasets. The results are presented in Table VII, with the FP32 precision results for Parallel Adapter directly taken from Table VI. Experimental results show that fine-tuning with low-precision backbone LLM does not significantly degrade model performance and they still maintain competitive accuracy. Our previous experiments have demonstrated that quantization can significantly reduce memory footprint during fine-tuning, making it suitable for resource-constrained edge environments.
VI-G Analysis the Benefits and Scalability of the Hybrid Parallelism Architecture.
@@ VI-F Analysis of Fine-Tuning with Quantized LLMs
Compared to standalone data parallelism or pipeline parallelism, the hybrid parallelism architecture adopted by PAC+ integrates the strengths of both DP and PP, enabling superior performance and scalability in resource-constrained edge environments. To explore the scalability advantages of PAC+’s hybrid parallelism over DP and PP, we compared the throughput of these methods when fine-tuning collaboratively across 2 to 8 Jetson Nano-H devices. The batch size was consistent with the number of devices, and the sequence length of each sample was fixed at 128. We implement DP and PP using the Parallel Adapters technique to ensure a fair comparison. Note that none of the three methods utilizes activation cache.
@@
Figure 18: Fine-tuning time with PAC+. Time without activation cache is represented by bars. The corresponding reduction in time achieved utilizing activation cache is represented by shaded areas. Dataset: MRPC.
Figure 16(b) illustrates the maximum per device memory footprint of model weights across edge cluster. For DP, each device must host a complete LLM, preventing the reduction of the parameters’ memory footprint through scaling up the number of devices. Therefore, as shown in Figure 16(a), the DP method exhibits OOM errors with both the BART-Large and T5-Large models. Conversely, PAC+ and PP partition the model into multiple stages with each handled by different devices. This approach allows for scaling the number of devices to reduce the peak memory footprint. PAC+’s hybrid parallelism offers a broader search space for parallel strategies compared to standalone PP. Our planning algorithm for PAC+ is capable of identifying more efficient hybrid parallel configurations within memory constraints, enhancing resource utilization. Although PAC+ may incur higher memory overhead in some instances, it achieves greater system throughput. Specifically, when compared to PP, PAC+ exhibits an increase in throughput from $39.50\%$ to $84.79\%$.
@@
To more clearly illustrate the parallel strategies adopted by PAC+, we present the device grouping configurations for PAC+ across various LLMs and numbers of devices in Figure 17. On the left, a table displays all the grouping results across three models. On the right, an instance is shown where a model is divided into three stages, with two devices handling the first stage to perform data parallelism. Specifically, when fine-tuning BART-Large with eight devices, DP encounters OOM issues because a single Jetson Nano cannot accommodate a complete BART-Large. PP addresses this problem by dividing the model into eight stages and employing straight pipeline parallelism for training. On the contrary, our PAC+ approach divides BART-Large into two stage models, with each stage replicated across four devices. This configuration significantly reduces the number of stages in the pipeline, thereby minimizing inter-stage data dependencies and communication latency, which in turn enhances the pipeline’s concurrent efficiency. These results demonstrate that our hybrid parallel approach offers a larger search space for parallel configurations, providing enhanced scalability and robustness across varying numbers of devices and workloads.
VI-H Comparison of PAC+ with and without Activation Cache.
@@
We further investigated how our activation cache mechanism benefits the required fine-tuning latency. By leveraging activation cache, the fine-tuning latency per epoch can decrease up to $79.51\%$. Figure 18 shows the fine-tuning latency reduction as the number of epochs increases with the use of the activation cache mechanism. We can observe that by leveraging the activation cache, fine-tuning latency is reduced by $26\%$ to $71\%$. Moreover, this reduction in latency increases with the number of epochs. For example, with the T5-Large model, training for two epochs reduces latency by $39\%$, whereas training for ten epochs increases the reduction to $71\%$. This reduction can be attributed to the fact that the Parallel Adapters constitute a lightweight, independent network, resulting in a significant decrease in training cost compared to the LLM backbone. We can bypass both the forward and backward passes through the LLM backbone since the necessary activations are already cached.
VII Related Work
@@
Parameter-Efficient Fine-Tuning for LLM. Prompt tuning [21] proposes to prepend the model input embeddings with a trainable tensor. Adapters tuning [13] adds domain-specific layers after attention and FFN layers in transformer. LoRA [16] decomposes the parameter update for a weight matrix into two trainable low-rank matrices. To further reduce the memory overhead, pioneering studies explore fine-tuning techniques that obviate the need for backpropagation through the backbone model. Y-tuning [27] learns additional task-specific label representations, which are integrated with the output of the backbone model to circumvent backpropagation. LST [37] involves the use of pruned lightweight transformer structures from the backbone as a side network. E3VA [48] extends the concept of the side network into the realm of computer vision.
Model Compressing Techniques for LLM Fine-tuning. Recent pioneering techniques focus on compressing LLMs to mitigate resource challenges. For instance, structural pruning, exemplified by LLM-Pruner [28], selectively removes non-critical coupled structures based on gradient information, thereby maximally preserving the majority of the LLM’s functionality. Similarly, researchers at Google [14] have demonstrated transferring knowledge from a large billion-parameter T5 teacher model to a student model with only millions of parameters via knowledge distillation, while retaining near-comparable performance. By reducing the model’s volume through compression, computational and memory resources required for full parameter fine-tuning can also be decreased.
The Parallel Adapter approach utilized in our work is categorized as a PEFT technique. However, our design is orthogonal to aforementioned model compression methods. Specifically, when the resource demands for the LLM backbone are substantial, the backbone can initially undergo appropriate compression to fit within the constraints of the target edge environment. Then, our Parallel Adapter technology can be further applied to significantly reduce the resources needed for fine-tuning, while simultaneously providing superior pluggable characteristics and mitigating catastrophic forgetting—critical advantages over full fine-tuning.
@@
On-device DNN Fine-Tuning. POET [34] achieves the finetuning of a BERT model on embedded devices, optimizing for both training speed and energy consumption. Lin et al. [26] enable training directly on devices with a minimal memory requirement of only 256KB. Sage and Melon [7, 40] implement hybrid memory management and conservation strategies, including operator fusion and the use of a dedicated memory pool, to mitigate memory limitations. Additionally, Mandheling [42] incorporates mixed-precision training along with DSP offloading to enhance the speed of learning.
@@
Collaborative Edge Computing for DNN Fine-Tuning. Federated Learning (FL) has been a promising paradigm in distributed machine learning that enables in-situ model fine-tuning. FwdLLM [44] designs a backpropagation-free fine-tuning FL protocol to enhance efficiency. AdaFL [2] proposes an FL framework for fine-tuning LLMs that features adaptable depth and width in its adapters modules. Breaking through the conventional paradigm of FL, Ye et al. [47, 50] devise a pipeline parallel architecture that facilitates the collaborative fine-tuning of DNNs across multiple edge devices. EDDL [11] adopts data parallelism training across embedded devices in a local area network. Asteroid [46] also employs HPP across multiple edge devices for DNN training, but it does not specifically address the parameter-efficient fine-tuning of LLMs.
This paper is an extension of the shorter conference version [32]. In this work, we further explore low-precision quantization of the LLM backbone and weight initialization for Parallel Adapters, enabling further optimization of both memory and time overhead in LLM fine-tuning. Also, we design a novel heterogeneity-aware parallelism planning algorithm to fully exploit the computational resource of heterogeneous edge clusters. Additionally, we provide a more detailed discussion of the design and implementation of our Parallel Adapters architecture.
VIII Conclusion
@@
This paper proposes PAC+, a time and memory efficient collaborative edge AI framework for personal LLMs fine-tuning. PAC+ breaks the resource wall of personal LLMs fine-tuning with a sophisticated algorithm-system co-design. Extensive evaluation of the prototype implementation demonstrates that PAC+ outperforms existing collaborative edge training systems, achieving up to a $9.7\times$ end-to-end speedup and reducing memory footprint by up to $88.16\%$ compared to mainstream LLM fine-tuning algorithms.